what is the slope perpindicular to 7x-3y=6?
Found 2 solutions by solver91311, checkley77:
Answer by solver91311(24713)



Step 1: Put the given equation into slope-intercept form by solving the equation for in terms of .

Step 2: Determine the slope of the given equation by inspection of the coefficient on in the slope-intercept form derived in Step 1.

Step 3: Use



to find the slope of the perpendicular. Which is to say, compute the negative reciprocal of the value discovered in Step 2.


John


Answer by checkley77(12844)

7x-3y=6?
-3y=-7x+6
y=-7x/-3+6/-3
y=7x/3-2 the slope of this line=7/3.
Therefore a perpendicular line would have a slope=-3/7(the negative reciprical).